OpenStudy (anonymous):

Evaluate each function for x = 2 and x = 7. h(x) = -4x + 61

OpenStudy (anonymous):

Can you put x = 2 in the equation??

OpenStudy (anonymous):

You have: h(x) = -4x + 61 Here if you put x = 2 then you will get h(2): \[h(2) = -4(2) + 61\] Can you solve this??

OpenStudy (anonymous):

@angel4ever35 I am expecting a reply from you side..

OpenStudy (anonymous):

106 or 52? Because once you first multiply then add, it's 52, so would that be the answer or 52(2)?

OpenStudy (anonymous):

53*

OpenStudy (anonymous):

See, 52 must be.. \[h(2) = -8 + 61\] \[h(2) = 53\] You will write like this only.. Now Second question: Put x = 7 as we did for x = 2, now put x = 7..

OpenStudy (anonymous):

5(7) = -28 + 61, 5(7) = 33

OpenStudy (anonymous):

Why are writing 5 in place of h??

OpenStudy (anonymous):

sorry I meant to hit h, It was an accident.

OpenStudy (anonymous):

h(7) = -28 + 61, 5(7) = 33*

OpenStudy (anonymous):

Again 5...

OpenStudy (anonymous):

I see h.

OpenStudy (anonymous):

opps h(7) = 33*

OpenStudy (anonymous):

ugh I don't understand why 5 keeps showing up.

OpenStudy (anonymous):

Yep.. Just concentrate and you will be on top..

OpenStudy (anonymous):

Well, thank you for your help. I'm not to great at math, haha.

OpenStudy (anonymous):

Who is great?? Just practice and concentrate, then you can definitely say I am Great at Math..

OpenStudy (anonymous):

Yea, I guess your right. Thank you again for your help. :)

OpenStudy (anonymous):

You are welcome dear..